# Present address: ]TRIUMF, Vancouver, BC V6T 2A3, Canada

### v1

@@ II Experimental Setup and Analysis
Nuclide ($X$) $I^{\pi}$ $T_{1/2}$ $R$ MEME [keV] $E_{x}$ [keV] MEMEa) $E_{x}$a) MEME b) $E_{x}$ b)
129gSn $3/2^{+}$ 2.23(4) m 0.969 963 694(20) -80 593.2 (25) -80 591(17)
129mSn $11/2^{-}$ 6.9(1) m 0.969 963 983(20) -80 557.4 (25) 35.8(35) -80 556(17) 35.15(5)
131gSn $3/2^{+}$ 56.0(5) s 0.985 038 825(16) -77 278.2 (19) -77 265(4)
131mSn $11/2^{-}$ 58.4(5) s 0.985 039 356(15) -77 212.4 (18) 65.8(26) -77 200(4) 65.1(3)
132gSb $(4)^{+}$ 2.79(7) m 0.992 543 964 0(82) -79 635.1 (10) -79 635.3(25) -79 629.6(27)
132mSb $(8^{-})$ 4.10(5) m 0.992 545 137 8(88) -79 489.8 (11) 145.3(15) -79 490(50)#\# 150(50)#\# -79 490.3(33) 139.3(20)
a)AME2020 and NUBASE2020 [AME20, NUBASE20].; b)Jaries et al[Jaries24].
@@ III Results
Our new results for 131g,mSn are a factor of 2 more precise than the AME2020 values [AME20]. Our values are approximately $3\sigma$ lighter than the AME value, which comes primarily from two previous TOF-ICR Penning trap measurements. Dworschak et al. [Dworschak08] could not resolve the ground and isomeric states, while Van Schelt et al. [VanSchelt13, VanSchelt12] did. Our results agree with Van Schelt et al. for the isomeric state very well, and our measured excitation energy is in good agreement with the value from NUBASE2020[NUBASE20]. We note that in production from spontaneously-fissioning 252Cf, the isomeric state is more populated than the ground state.

### v2

@@ II Experimental Setup and Analysis
Nuclide ($X$) $I^{\pi}$ $T_{1/2}$ [m] $R$ MEME [keV] $E_{x}$ [keV] MEMEa) $E_{x}$a) MEME $E_{x}$
129gSn $3/2^{+}$ 2.23(4) 0.969 963 694(20) -80 593.2 (25) -80 591(17)
129mSn $11/2^{-}$ 6.9(1) 0.969 963 983(20) -80 557.4 (25) 35.8(35) -80 556(17) 35.15(5)
131gSn $3/2^{+}$ 0.933(8) 0.985 038 825(16) -77 278.2 (19) -77 265(4)
131mSn $11/2^{-}$ 0.973(8) 0.985 039 356(15) -77 212.4 (18) 65.8(26) -77 200(4) 65.1(3) 64(2)b)
132gSb $(4)^{+}$ 2.79(7) 0.992 543 964 0(82) -79 635.1 (10) -79 635.3(25) -79 629.6(27)c)
132mSb $(8^{-})$ 4.10(5) 0.992 545 137 8(88) -79 489.8 (11) 145.3(15) -79 490(50)#\# 150(50)#\# -79 490.3(33)c) 139.3(20)c)
a)AME2020 and NUBASE2020 [AME20, NUBASE20].; b)Benito et al[Benito24]. c)Jaries et al[Jaries24].
@@ III Results
Our new results for 131g,mSn are a factor of 2 more precise than the AME2020 values [AME20]. Our values are approximately $3\sigma$ lighter than the AME value, which comes primarily from two previous TOF-ICR Penning trap measurements. Dworschak et al. [Dworschak08] could not resolve the ground and isomeric states, while Van Schelt et al. [VanSchelt13, VanSchelt12] did. Our results agree with Van Schelt et al. for the isomeric state very well, and our measured excitation energy is in good agreement with the value from NUBASE2020[NUBASE20] as well as the excitation energy directly measured using the PI-ICR technique at JYFLTRAP [Benito24]. We note that in production from spontaneously-fissioning 252Cf, the isomeric state is more populated than the ground state.